A hummingbird hovers by exerting a downward force on the air equal, on average, to its weight. By Newton's third law, the air exerts an upward force of the same magnitude on the bird's wings. Find the average mechanical power (in W) delivered by a 3.40 g hummingbird while hovering if its wings beat 55.0 times per second through a stroke 3.34 cm long.


__W


1
Expert's answer
2021-09-23T08:31:20-0400

"Force\; applied = m \times g

= \frac{3.4}{1000} \times 9.8 \\

= 33.32 \times 10^{-3} \;N"

Strike length "= 3.34 \; cm = 3.34 \times 10^{-2} \;m"

Work done = "Force \times distances"

"= 33.32 \times 10^{-3} \times 3.34 \times 10^{-2} \\

= 111.288 \times 10^{-5} \;J"

Frequency of bead = 55.0 /s

Time of bead "= \frac{1}{55}"

"Power = \frac{work}{time} \\

= \frac{111.288 \times 10^{-5}}{1/55} \\

= 0.061208 \;J/s"

The average mechanical power = 0.061208 W